Question 1 Evaluate 5 23-3 (28 – 1).
Answer #1

S 6 -5 < s=3 Given : Therefore applying linit (2(3)-1) + (2 (4) -1) + (215)-1) » -(6-1)+ (8-1) + (10-1) I . T = ) S + 7 + 9 =please rate my answer with thumbs up if it helps